# Simulating Relational Event Histories - Why and How

### v2

@@ Abstract
Many important social phenomena result from repeated interactions among individuals over time such as email exchanges in an organization, or face-to-face interactions in a classroom. Insights into the mechanisms underlying the dynamics of these interactions can be achieved through simulations of networks on a fine temporal granularity. In this paper, we present statistical frameworks to simulate relational event networks under dyadic and actor-oriented relational event models. These simulators have a broad applicability in temporal social network research such as model fit assessment, theory building, network intervention planning, making predictions, understanding the impact of network structures, to name a few. We show this in three extensive applications. First, it is shown why simulation-based techniques are crucial for relational event model assessment, for example to investigate how past events affect future interactions in the network. Second, we show how simulation techniques are important when building and extending theories about social phenomena such as understanding social identity dynamics using optimal distinctiveness theory. Third, we demonstrate how simulation techniques contribute to a better understanding of the longevity of network interventions.
@@ 1 Introduction
Our understanding of social interaction mechanisms between individuals has been enhanced by dynamic network-based approaches. A network approach allows for the representation of individuals by vertices and an event or interaction between the individuals by the edges between the vertices. In traditional, static network analysis, edges are assumed to be stable and their presence or absence is indicative for the state the network is in. In dynamic network analysis, edges (and vertices, for that matter) are allowed to appear and disappear. In most dynamic network applications, one observes a network repeatedly and then notes for each edge its presence or absence at observation time. In continuous time dynamic network analysis (of which this paper is an example), edges are observed to appear and dissolve in real time at any time. To distinguish between edges that are stable and edges that can be in constant flux, the latter are often termed “relational events.” The measurement of network changes over time gives access to fine-grained temporal information about the evolution of the network. This allows for investigation of the relationship between individual behaviour of individual nodes or dyads in the network, and collective behaviour of the network. Particularly the temporal information yields insight into how variation in individual level dynamics affect network level characteristics, which in turn further affect individual behaviour. Several statistical methods have been developed to study dynamic relational event or longitudinal social network data (Holland and Leinhardt,, 1977; Snijders et al.,, 2010; Hanneke et al.,, 2010; Krivitsky and Handcock,, 2014; Perry and Wolfe,, 2013). Of particular interest to us is the relational event model introduced by Butts, (2008), and Stadtfeld, (2014) which allows for modelling of dynamic networks when the exact time-stamp of the events are available. Relational event models are a class of models that model the event rate in a dynamic network as a function of the network’s past history and exogenous information available about the actors (nodal covariates) or pairs of actors (dyadic covariates) in the network. These models provide a flexible framework for network inference that allows for estimation of various network effects such as homophily, reciprocity, transitivity, and preferential attachment while jointly accounting for the fine-grained temporal nature of real-life social interaction.
Despite its usefulness to quantify the relative importance of various effects on the network evolution, it may still be difficult to truly understand the underlying social mechanisms based on a fitted relational event model. For example, it may be that several relational event models achieve similar fit to the data or the true underlying social mechanisms may be too complex to discover using a relational event model only. This is echoed in Box, (1976) famous quote that “all models are wrong…”. It is possible, however, to get a better understanding about the true underlying social mechanisms by further exploring what network structures look like under different relational event modeling settings via simulations. Simulations can provide an approximation of applied research settings as long as the simulations generate realistic networks. A major feature of the relational event-based approach to simulations, making it a suitable choice for simulating realistic datasets, is its flexibility in allowing for a wide range of network effects and covariates that can be utilized to generate datasets with many complex mechanisms in juxtaposition.
In this paper we introduce general frameworks to simulate relational event networks under the dyadic (Butts,, 2008) and the actor oriented (Stadtfeld and Block,, 2017) relational event model that can be utilized for simulating relational event sequences. The simulation frameworks can be used for various objectives:
@@
This article has several goals. In Section 2, we introduce the dyadic and actor oriented relational event models and outline the general-purpose simulation frameworks for these models that can be customized in various ways using problem-specific information. Next we illustrate a range of simulation scenarios that showcase the broad potential of the simulation frameworks in an attempt to inspire and guide new research. In Section 3 we demonstrate how the simulations can be used to assess goodness of fit of relational event models using an email network from an organization. In Section 4, we describe how simulations can be used to develop theories about social phenomena that drive interactions and provide an example by simulating a well known social theory and test a boundary condition on this theory using simulations. In Section 5 we discuss how the simulation framework can be used to investigate outcomes of network-based interventions. We demonstrate how strategies for network-based interventions can be simulated and compared on an organizational network. Finally, we conclude with a discussion on limitations and future prospects of the simulation techniques in Section 6.
@@ 2.1 Simulation Framework 1: Relational Event Model (REM)
Consider a sequence of $M$ events $E=\{e_{1},e_{2},\dots e_{M}\}$ such that. $t_{m}<t_{m+1}$. An event $e_{m}=(i_{m},j_{m},t_{m})$, entails a sender $i_{m}\in\mathcal{A}$, a receiver $j_{m}\in\mathcal{A}$, and the time at which the event was observed, $t_{m}$. The set of possible events that can occur at any time is referred to as the risk set $\mathcal{R}\subseteq\{(i,j):i,j\in\mathcal{A}\}$. The risk set may be fixed or vary with time endogenously, i.e particular events may give rise to the possibility of other events that were previously impossible, or the risk set may vary exogenously, where the possibility of occurrence of certain events is determined by exogenous factors such as availability of actors, location, or other circumstances specific to the research setting  [^3] Without loss of generality, we will keep the riskset constant in this paper..
In REM, each dyad has its own rate of occurrence. A higher rate implies that the event involving that dyad is more likely to occur soon, and a lower rate implies that the occurrence of the event is more rare. The rate is modeled as a log linear function of the endogenous and exogenous statistics pertaining to that dyad along with the parameters $\beta$ that represent the strength of these statistics to explain social interaction behavior in the network. The occurrence of events is modelled using a piece-wise constant hazard model. Under the piece-wise constant model, the rates are assumed to only change when an event occurs (anywhere in the network). Thus, as time progresses and events are observed, the rate is updated to reflect the new network history. The rate $\lambda^{dyadic}_{ij}$ of a dyad $(i,j)$ at time $t$ is then specified in a log linear form as:
$\lambda^{dyad}_{ij}(t)=\begin{cases}\text{exp}\{\beta^{T}\;X(i,j,E_{t})\}&\ (i,j)\in\mathcal{R}\\ 0&\ (i,j)\notin\mathcal{R}\end{cases}$ (1)
where $X(i,j,E_{t})$ is a vector of $P$ statistics for the dyad $(i,j)$ on the event sequence $E_{t}$ until time t and $\beta\in{\rm I\!R^{P}}$ is the vector of corresponding parameters. The statistics vector $X(i,j,E_{t})$ can capture either endogenous network characteristics as a function of the past events in the history $E_{t}$ before time $t$ or exogenous actor or dyadic attributes. The notation for rate has a superscript ‘dyad’ to distinguish it from the rate under the actor-oriented model introduced in the next section.
@@
$p^{dyad}(\delta_{m}|E_{t_{m-1}},\beta,X)\sim\text{Exponential}(\Lambda^{dyad}(t_{m-1}))$ (2)
@@
$\begin{split}p^{dyad}(i,j\mid t,\beta,X,E_{t})=\frac{\lambda^{dyad}_{ij}(t)}{\Lambda^{dyad}(t)}.\;\end{split}$ (3)
In order to simulate relational event networks, the user needs to specify which endogenous or exogenous network statistics (e.g., Leenders et al.,, 2016) are included in the model and the magnitude of the corresponding parameters (based on theory, a specific dynamic of theoretical interest, or a fitted model). If preferred, it also possible to initialize simulations with a pre-defined starting sequence $E_{0}$. This would be useful if a user wants to predict the immediate future of an empirical network or to simulate using different models from the same event history. The inputs to the simulation algorithm play an important role in determining the characteristics of the sequence output. The number of actors $N$ in the network determines the size of the riskset as $N(N-1)$ (when all dyads are at risk) and $\tau$, the time until which the simulation algorithm is run, determines the number of events in the sequence. The parameters $\beta$ capture the sign and magnitude of the effect of the statistics. Given these inputs, a relational sequence $E$ can be simulated using Algorithm 1.
input : $\mathcal{\beta},\tau,\mathcal{R}$, $E_{o}$
@@ 4      Initialize event sequence $E=E_{o}$
9      update event rates $\lambda^{dyad}_{ij}(t_{m-1}\mid E_{t_{m-1}},\beta,X)\ \forall(i,j)\in\mathcal{R}$ using Equation (1)
@@ 2.2 Simulation Framework 2: Dynamic Network Actor Model (DyNAM)
$p^{actor}(\delta_{m}|E_{t},\beta,X^{s})\sim\text{Exponential}(\Lambda^{sender}(t)),$
@@
$\lambda_{i}^{sender}(t)=\exp\{\gamma^{T}X^{s}(i,E_{t}),\}$ (4)
@@
$p^{actor}(i\mid E_{t},\gamma,X^{s})=\frac{\lambda_{i}^{sender}(t)}{\Lambda^{sender}(t).}$ (5)
@@
$p^{actor}(j|i,t,E_{t},\alpha,X^{r})=\frac{\lambda_{j|i}^{receiver}(t)}{\Lambda_{i}^{receiver}(t)}$ (6)
@@
$\lambda_{j|i}^{receiver}(t)=\text{exp}(\alpha^{T}X^{r}_{ij}),$
@@
input : $\gamma,\alpha,\tau,\mathcal{R},E_{o}$
@@ 5      Initialize event sequence $E=E_{o}$
9       compute sender rates $\lambda_{i}(t_{m-1}\mid E_{t_{m-1}},\gamma,X^{s})\ \forall i\in\mathcal{A}$ using Equation (4)
@@
15       update $\beta$ if time-varying
@@ 3 Assessing goodness-of-fit through simulations
In the following we will present three extensive applications that highlight why relational event simulators are essential for temporal social network research. We begin with assessing goodness of model fit through simulations.
@@
Typical approaches to evaluating goodness-of-fit of relational event models involve a balance of parsimony and accuracy, i.e. choosing a model with fewest parameters and with satisfactory accuracy for predicting events. This usually involves using information criteria such as the Akaike Information Criterion (AIC), the Bayesian Information Criterion (BIC), or other likelihood-based goodness-of-fit measures. These criteria are useful as a relative measure of fit when comparing models but they do not give an indication of the model fitness in an absolute sense neither do they provide a substantive direction of misfit. For example, when comparing multiple models using general information criteria, the “best” model may still have a poor fit to the data in terms of capturing important network characteristics, however, and it may result in poor predictions. On the other hand, simulation-based methods can be used to assess model fit in an absolute sense by assessing whether important network characteristics in the data are also present in simulated data using the fitted model. If simulated sequences based on a fitted model bear little resemblance to the observed sequence (or misses important characteristics of the network dynamic), this generally suggests a poor fit of the model to the data. Misfit can occur due to inclusion of effects that are not operative in the data or due to exclusion of key effects or external factors that play an important role in the network dynamics. Inferences based on misfitted models are therefore (theoretically) unreliable, and simulation-based methods are needed to assess that the model results in a reasonable fit to the observed data sequence before drawing inferences.
@@
Simulation-based methods for model assessment have also been extensively used in other directions of social network research. Hunter et al., (2008) introduced simulation-based goodness-of-fit tests for social networks under exponential random graph models (ERGMs) using structural indices such as degree distributions, edgewise shared partners, and geodesic distance distributions. Snijders and Steglich, (2015) introduce a variety of structural fit indices to describe cross sectional data such as the size of the largest component, number of components, median geodesic distance, transitivity coefficient, variance of in or out degree divided by mean, correlation between in and out degrees, graph hierarchy and, least upperboundedness. Lospinoso and Snijders, (2019) include higher order indices in the form of triad census to assess if the simulated networks accurately resemble triadic structures in the observed network. Wang et al., (2022) describe behavioral indices to assess fit for models that jointly model behaviour and network structure. Brandenberger, (2019) considered the accuracy of event predictions to assess the goodness-of-fit of relational event models for political networks. Because relational event models not only model the network structure but the timing and order of events as well, goodness-of-fit tests would be incomplete without assessing the fit of candidate models on temporal indices. It is possible that the model fits data well within certain time periods but not in others or that events between certain dyads in the predicted dataset are further apart than in the observed network. In addition, it is possible to compute structural indices that respect the timing and order of events in relational event datasets, such as temporal degree, temporal reachability, and temporal betweenness (Falzon et al.,, 2018). Nicosia et al., (2013) similarly define temporal connectedness, various temporal centrality indices and temporal shortest paths for time-respecting paths in dynamic networks. These temporal indices enhance the fit indices used for static or longitudinal networks, by also incorporating the timing information available from relational event datasets. The measures mentioned above are just some of the fit indices used in social network literature; there are abundant other measures that can also be considered. It is impossible to define an exhaustive set of characteristics for general use with the relational event model. Rather than searching for the best fitting model, we believe researchers should focus on a model that reproduces specific aspects of the data that are of theoretical or practical interest rather than applying a default set of statistics. Ultimately, the choice of the fit indices should depend on the specific phenomena the researcher intends to address in their analysis.

### v4

# Simulating Relational Event History Data: Why and How

@@ Abstract
Many important social phenomena are characterized by repeated interactions among individuals over time such as email exchanges in an organization or face-to-face interactions in a classroom. To understand the underlying mechanisms of social interaction dynamics, statistical simulation techniques of longitudinal network data on a fine temporal granularity are crucially important. This paper makes two contributions to the field. First, we present statistical frameworks to simulate relational event networks under dyadic and actor-oriented relational event models which are implemented in a new R package remulate. Second, we explain how the simulation framework can be used to address challenging problems in temporal social network analysis, such as model fit assessment, theory building, network intervention planning, making predictions, understanding the impact of network structures, to name a few. This is shown in three extensive case studies. In the first study, it is elaborated why simulation-based techniques are crucial for relational event model assessment which is illustrated for a network of criminal gangs. In the second study, it is shown how simulation techniques are important when building and extending theories about social phenomena which is illustrated via optimal distinctiveness theory. In the third study, we demonstrate how simulation techniques contribute to a better understanding of the longevity and the potential effect sizes of network interventions. Through these case studies and software, researchers will be able to better understand social interaction dynamics using relational event data from real-life networks.
Author profiles: Rumana Lakdawala: https://scholar.google.com/citations?hl=en&user=x4z23RAAAAAJ Joris Mulder: https://scholar.google.com/citations?hl=en&user=W8EG7_gAAAAJ Roger Leenders: https://scholar.google.com/citations?hl=en&user=63GISngAAAAJ
Keywords: relational events; simulation techniques; temporal social networks; model fit assessment; interventions; actor-oriented models; dyadic interaction models
@@ 1 Introduction
Our understanding of social interaction mechanisms between actors has been enhanced by dynamic network-based approaches. These actors can represent individuals, groups of individuals, organizations, or even countries. A network approach allows for the representation of actors by vertices and an event or interaction between the actors by the edges between the vertices. In traditional, static network analysis, edges are assumed to be stable and their presence or absence is indicative for the state the network is in. In continuous time dynamic network analysis (which is the focus of this paper), edges can appear and/or dissolve in real time at any time. To distinguish between edges that are stable and edges that can be in constant flux, the latter are often termed “relational events.” The analysis of relational event data has the potential for a deeper understanding of the underlying social structures and how these evolve over time. This allows for the investigation of relationships between actors, dyads, and collective behaviour in the network. Particularly the temporal information yields insight into how variation in individual level dynamics affects network level characteristics, which in turn further affect individual behaviour. Several statistical methods have been developed to study dynamic relational event history data for longitudinal social network analysis (Holland and Leinhardt,, 1977; Snijders et al.,, 2010; Hanneke et al.,, 2010; Krivitsky and Handcock,, 2014; Perry and Wolfe,, 2013). Of particular interest to us is the relational event models proposed by Butts, (2008) and Stadtfeld, (2014), which allow dynamic social network analyses when the exact time-stamp of the events are available. In these relational event models, the event rate between actors is modeled as a function of the actors’ past history and exogenous information available about the actors (nodal covariates) or pairs of actors (dyadic covariates) in the network. These models provide a flexible framework for network inference that allows for estimation of various drivers of social interaction behavior, such as homophily, reciprocity, transitivity, and preferential attachment.
Despite its usefulness to quantify the relative importance of network drivers of the social evolution, truly understanding the underlying social mechanisms between actors based on relational event history data remains an enormous challenge. For example, it may be that different relational event models may induce similar social interaction behavior or the true underlying social mechanisms may be too complex to discover using a relational event analysis only. This is echoed in Box, (1976) famous quote that “all models are wrong…”. Simulation techniques have the potential to tackle these important current challenges in temporal social network modeling due to their ability to assess the impact of model specifications on data characteristics in a direct manner.
For this reason, the aims of the current paper is two fold. First it further develops and extends simulation techniques under flexible relational event modeling frameworks. Simulation frameworks are proposed under dyadic relational event models (Butts,, 2008) and under actor-oriented relational event models (Stadtfeld and Block,, 2017). Unlike existing approaches, the proposed simulators support time-varying network effects, constrained risk sets, and various memory decay function to compute endogenous statistics allowing the study of the elapsed time of past events on the event rate in the (near) future. Second, the paper explores the potential of simulation techniques to improve social network analyses and to better understand social interaction behavior in real-life applications. Specifically, relational event simulators can contribute to the following objectives:
@@
To keep the paper as concise as possible, the paper shall focus on the first three objectives: goodness of fit assessment of relational event models, theory building on social interaction dynamics, and the planning network interventions.
This article is organized as follows. In Section 2, we introduce the dyadic and actor-oriented relational event models and outline the general-purpose simulation frameworks for these models that can be customized in various ways using problem-specific information. Next we illustrate a range of simulation scenarios that showcase the broad potential of the simulation frameworks in an attempt to inspire and guide new research. In Section 3 we demonstrate how the simulations can be used to assess goodness of fit of relational event models using an email network from an organization. In Section 4, we describe how simulations can be used to develop theories about social phenomena that drive interactions and provide an example by simulating a well known social theory and test a boundary condition on this theory using simulations. In Section 5 we discuss how the simulation framework can be used to investigate outcomes of network-based interventions. We demonstrate how strategies for network-based interventions can be simulated and compared on an organizational network. Finally, we conclude with a discussion on limitations and future prospects of the simulation techniques in Section 6.
@@ 2.1 Simulation Framework 1: Relational Event Model (REM)
Consider a sequence of $M$ events $E=\{e_{1},e_{2},\dots e_{M}\}$ such that. $t_{m}<t_{m+1}$. An event $e_{m}=(i_{m},j_{m},t_{m})$, entails a sender $i_{m}\in\mathcal{A}$, a receiver $j_{m}\in\mathcal{A}$, and the time at which the event was observed, $t_{m}$. The set of possible events that can occur at any time is referred to as the risk set $\mathcal{R}\subseteq\{(i,j):i,j\in\mathcal{A}\}$. The risk set may be constrained such that certain actors cannot interact with one another. Therefore the constrained riskset can be fixed or vary exogenously, where the possibility of occurrence of certain events is determined by exogenous factors such as availability of actors, location, or other circumstances specific to the research setting.
In REM, each dyad has its own rate of occurrence. A higher rate implies that the event involving that dyad is more likely to occur soon, and a lower rate implies that the occurrence of the event is more rare. The rate is modeled as a log linear function of the endogenous and exogenous statistics pertaining to that dyad along with the parameters $\boldsymbol{\beta}(t)$ that represent the strength of these statistics to explain social interaction behavior in the network. The occurrence of events is modelled using a piece-wise constant hazard model. Under the piece-wise constant model, the rates are assumed to only change when an event occurs (anywhere in the network). Thus, as time progresses and events are observed, the rate is updated to reflect the new network history. The rate $\lambda^{dyadic}_{ij}$ of a dyad $(i,j)$ at time $t$ is then specified in a log linear form as:
$\lambda^{dyad}_{ij}(t)=\begin{cases}\text{exp}\{\boldsymbol{\beta}(t)^{T}\;X(i,j,E_{t})\}&\ (i,j)\in\mathcal{R}\\ 0&\ (i,j)\notin\mathcal{R}\end{cases}$ (1)
where $X(i,j,E_{t})$ is a vector of $P$ statistics for the dyad $(i,j)$ on the event sequence $E_{t}$ until time t and $\boldsymbol{\beta}(t)\in{\rm I\!R^{P}}$ is the vector of corresponding parameters at time $t$ (the parameters may be constant and not-time varying in which case, $\beta$ is a constant. The statistics vector $X(i,j,E_{t})$ can capture either endogenous network characteristics as a function of the past events in the history $E_{t}$ before time $t$ or exogenous actor or dyadic attributes. The notation for rate has a superscript ‘dyad’ to distinguish it from the rate under the actor-oriented model introduced in the next section.
@@
$p^{dyad}(\delta_{m}|E_{t_{m-1}},\boldsymbol{\beta}(t_{m-1}),X)\sim\text{Exp}(\Lambda^{dyad}(t_{m-1}))$ (2)
@@
$\begin{split}p^{dyad}(i,j\mid t,\boldsymbol{\beta}(t),X,E_{t})=\frac{\lambda^{dyad}_{ij}(t)}{\Lambda^{dyad}(t)}.\;\end{split}$ (3)
In order to simulate relational event networks, the user needs to specify which endogenous or exogenous network statistics (e.g., Leenders et al.,, 2016) are included in the model and the magnitude of the corresponding parameters (based on theory, a specific dynamic of theoretical interest, or a fitted model). If preferred, it also possible to initialize simulations with a pre-defined starting sequence $E_{0}$. This would be useful if a user wants to predict the immediate future of an empirical network or to simulate using different models from the same event history. The inputs to the simulation algorithm play an important role in determining the characteristics of the sequence output. The number of actors $N$ in the network determines the size of the riskset as $N(N-1)$ (when all dyads are at risk) and $\tau$, the time until which the simulation algorithm is run, determines the number of events in the sequence. The parameters $\boldsymbol{\beta}(t)$ capture the sign and magnitude of the effect of the statistics. Given these inputs, a relational sequence can be simulated using Algorithm 1.
input : $\boldsymbol{\beta},\tau,\mathcal{R}$, $E_{o}$
@@ 4      Initialize event sequence $E=E_{o}$
9      update event rates $\lambda^{dyad}_{ij}(t_{m-1}\mid E_{t_{m-1}},\boldsymbol{\beta}(t),X)\ \forall(i,j)\in\mathcal{R}$ using Equation (1)
@@ 2.2 Simulation Framework 2: Dynamic Network Actor Model (DyNAM)
$p^{actor}(\delta_{m}|E_{t},\boldsymbol{\gamma}(t),X^{s})\sim\text{Exp}(\Lambda^{sender}(t)),$
@@
$\lambda_{i}^{sender}(t)=\exp\{\boldsymbol{\gamma}(t)^{T}X^{s}(i,E_{t}),\}$ (4)
@@
$p^{actor}(i\mid E_{t},\boldsymbol{\gamma}(t),X^{s})=\frac{\lambda_{i}^{sender}(t)}{\Lambda^{sender}(t).}$ (5)
@@
$p^{actor}(j|i,t,E_{t},\boldsymbol{\alpha}(t),X^{r})=\frac{\lambda_{j|i}^{receiver}(t)}{\Lambda_{i}^{receiver}(t)}$ (6)
@@
$\lambda_{j|i}^{receiver}(t)=\text{exp}(\boldsymbol{\alpha}(t)^{T}X^{r}(i,j,E_{t}),$
@@
input : $\boldsymbol{\gamma},\boldsymbol{\alpha},\tau,\mathcal{R},E_{o}$
@@ 5      Initialize event sequence $E=E_{o}$
9       compute sender rates $\lambda_{i}(t_{m-1}\mid E_{t_{m-1}},\boldsymbol{\gamma}(t),X^{s})\ \forall i\in\mathcal{A}$ using Equation (4)
@@
15       update $\boldsymbol{\gamma}(t)$ and $\boldsymbol{\alpha}(t)$ if time-varying
@@ 16       $m++$
2.3 remulate: R package to simulate relational event histories
This article introduces the R package remulate (Lakdawala et al.,, 2020). The package is developed to assist researchers in simulating relational event histories. The package aims to reduce the complexity associated with simulation techniques and to make them more accessible to social network researchers. The package enables the user to simulate using a wide range of commonly used exogenous and endogenous statistics for the tie- and actor-oriented relational event model approaches as well as several important extensions.
Unlike relevent (Butts,, 2023) which also allows users to simulate relational event histories, remulate has the following additional features:
  1. 1.
    Support for dyadic relational events models as well as actor-orientied relational event models.
  2. 2.
    Support of a larger collection of pre-computed endogenous statistics with various methods of normalization and standardization and support for interaction terms.
  3. 3.
    Support different types of memory decay functions, such as exponential decays or step-wise functions, to study the influence of the transpired time of past events on the event rate between actors in the future.
  4. 4.
    Support for constrained risk sets, such that if two actors cannot interact, events involving that dyad would not be sampled.
  5. 5.
    Support time-varying network parameters to study instantaneous, gradual, and periodic changes of drivers of social interaction behavior.
  6. 6.
    Support for frailty relational event models to capture nodal heterogeneity.
  7. 7.
    Support for relational event block models.
Appendix A presents exemplary R code on how to simulate from tie-oriented and actor-oriented relational event models using the remulate package.
@@ 3 Assessing goodness-of-fit through simulations
In the following we will present extensive applications that highlight why relational event simulators are essential for temporal social network research. We begin with assessing goodness of model fit through simulations.
@@
Typical approaches to evaluating goodness-of-fit of relational event models involves a comparison of different models by balancing parsimony and accuracy. The preferred model generally contains the fewest parameters while yielding satisfactory fit for the observed relational events. This usually involves using information criteria such as the Akaike Information Criterion (AIC), the Bayesian Information Criterion (BIC), or other likelihood-based goodness-of-fit measures. Although these criteria are useful as a relative measure of fit when comparing models, they do not give an indication of the model fitness in an absolute sense neither do they provide a substantive direction of possible misfit. Thus, when comparing multiple relational event models using such general information criteria, the “best” model may still have a very poor fit to the observed event sequence in terms of capturing important network characteristics and it may result in poor predictions.
On the other hand, simulation-based methods can be used to assess model fit in an absolute sense by assessing whether important network characteristics in the data are also present in the simulated data using the fitted model. If simulated sequences based on a fitted model bear little resemblance to the observed sequence (or misses important characteristics of the network dynamic), this generally suggests a poor fit of the model to the data. Misfit can occur due to inclusion of effects that are not operative in the data or due to exclusion of key effects or external factors that play an important role in the network dynamics. Inferences based on misfitted models are therefore (theoretically) unreliable, and simulation-based methods are needed to assess whether the model results in a reasonable fit to the observed data sequence before drawing inferences.
@@
Before describing the general methodology, we briefly mention some previous work on Simulation-based methods for model assessment in social network research. Hunter et al., (2008) introduced simulation-based goodness-of-fit tests for social networks under exponential random graph models (ERGMs) using structural indices such as degree distributions, edgewise shared partners, and geodesic distance distributions. Snijders and Steglich, (2015) introduce a variety of structural fit indices to describe cross sectional data such as the size of the largest component, number of components, median geodesic distance, transitivity coefficient, variance of in or out degree divided by mean, correlation between in and out degrees, graph hierarchy and, least upperboundedness. Lospinoso and Snijders, (2019) include higher order indices in the form of triad census to assess if the simulated networks accurately resemble triadic structures in the observed network. Wang et al., (2020) describe behavioral indices to assess fit for models that jointly model behaviour and network structure. Brandenberger, (2019) considered the accuracy of event predictions to assess the goodness-of-fit of relational event models for political networks. Because relational event models not only model the network structure but the timing and order of events as well, goodness-of-fit tests would be incomplete without assessing the fit of candidate models on temporal indices. It is possible that the model fits data well within certain time periods but not in others or that events between certain dyads in the predicted dataset are further apart than in the observed network. In addition, it is possible to compute structural indices that respect the timing and order of events in relational event datasets, such as temporal degree, temporal reachability, and temporal betweenness (Falzon et al.,, 2018). Nicosia et al., (2013) similarly define temporal connectedness, various temporal centrality indices and temporal shortest paths for time-respecting paths in dynamic networks. These temporal indices enhance the fit indices used for static or longitudinal networks, by also incorporating the timing information available from relational event datasets. In addition Amati et al., (2024) suggest examining internal time structures in relational event models to capture the specific sequence and timing of events, such as reciprocity and transitive closure to ensure that the model accurately reflects the temporal patterns observed in social interactions.
The indices mentioned above are just a few examples from the social network literature to assess the goodness of fit. In fact the number of temporal network characteristic to assess is unlimited. Moreover when assessing goodness of fit, we believe that researchers should focus on reproducing specific aspects of the data that are of theoretical and/or practical interest rather than applying a default set of statistics. Ultimately, the choice of the fit indices should depend on the specific phenomena the researchers intend to address in their analysis. Therefore, it is impossible to provide an exhaustive set of network characteristics for general use in relational event analysis. Instead, we provide a general step-by-step methodology to assess the goodness of fit in relational event modeling applications. Subsequently, we apply the methodology to a real-life case study.